OpenStudy (anonymous):

Which Value(s) of X are excluded from the solution to the following equation ? Show all of your work to get to your solution . x^2-5x-20 x 2 ----------------- = -------- + -------- x^2-17x+70 x-10 x-7

OpenStudy (anonymous):

I got this

OpenStudy (anonymous):

Do you know how to factorize a given polynomial??

OpenStudy (anonymous):

No, i dont know how to factorize a given polynomial. @waterineyes

OpenStudy (anonymous):

What will you put as value of x in the denominator so that you will get 0 in the denominator??

OpenStudy (anonymous):

\[\frac{ x^2-5x-20 }{ x^2-17x+70 }=\frac{ x }{ x-10 }+\frac{ 2 }{ x-7 }\]

OpenStudy (anonymous):

See, I think we need not to calculate x here, we just want that what values of x should be excluded so that we can have well defined values of x in this question.

OpenStudy (anonymous):

@Forever_Me , you know that you can't divide by zero?

OpenStudy (anonymous):

Yep.. Division by zero should be avoided, as we can't do this..

OpenStudy (anonymous):

find the values of x so that x-10=0 x-7=0 x^2-17x+70=0

OpenStudy (anonymous):

In RHS, we have: \[\frac{1}{x - 10}\] Can you tell us that for what value of x, denominator (x-10) will be zero??

OpenStudy (anonymous):

Yes , i know we cant divide by Zero @Peter14

OpenStudy (anonymous):

WOuld you mind explaning exactly what you want me to do . @waterineyes

OpenStudy (anonymous):

Just put (x-10) = 0 So you will get: \[x - 10 = 0\] So what will be x, can you tell?

OpenStudy (anonymous):

Hey, you have to just do the calculation. Add 10 both the sides and find x..

OpenStudy (anonymous):

So x= 10 ? @waterineyes